# Andreas Schlaginhaufen SYCAMORE, EPFL

### v1

@@ 5 Experiments
To validate our results experimentally, we adopt a stochastic variant of the WindyGridworld environment (Sutton and Barto, 2018). In this environment, the agent moves to the intended grid cell with a probability of $(1-\beta)$ and is pushed one step further in the direction of the wind with a probability of $\beta$. Using Algorithm 1, we recover a reward $\hat{r}$ from demonstrations of two experts, both exposed to the same wind strength $\beta$ but different wind directions – North and East. The experiments are repeated for a varying number of expert demonstrations $N^{\textsf{E}}\in\{10^{3},10^{4},10^{5},10^{6}\}$ and wind strengths $\beta\in\{0.01,0.1,0.5,1.0\}$. We then test the transferability to two different environments: one with South wind, $P^{\text{South}}$, and a zero-wind environment with cyclically shifted actions, $P^{\text{Shifted}}$. Figure 2(a) shows that the second principal angle between the two experts’ transition laws $P_{0}$ and $P_{1}$ increases with increasing wind strength. Moreover, Figure 2(b)-(d) show that both the closeness between $\hat{r}$ and $r^{\textsf{E}}$ in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$ and the transferability to $P^{\text{South}}$ and $P^{\text{Shifted}}$ improve with a larger second principal angle, as expected from Theorem 3.4. For a more detailed discussion of the experiments we refer to Appendix I.
(a)                         (b)                         (c)                         (d)
Figure 2: (a) shows the second principal angle between the experts, for varying wind strength $\beta$. (b) shows the distance between $\hat{r}$ and $r^{\textsf{E}}$ in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$ for a varying number of expert demonstrations $N^{\textsf{E}}$ and wind strength $\beta$. (c) and (d) show the transferability to $P^{\text{South}}$ and $P^{\text{Shifted}}$ in terms of $\ell_{P^{\text{South}}}(r^{\textsf{E}},\mathsf{RL}_{P^{\text{South}}}(\hat{r}))$ and $\ell_{P^{\text{Shifted}}}(r^{\textsf{E}},\mathsf{RL}_{P^{\text{Shifted}}}(\hat{r}))$, respectively. The circles indicate the median and the shaded areas the 0.2 and 0.8 quantiles over the 10 independent realizations.
@@ 6 Conclusion
Our results provide only sufficient conditions for transferability. It would be valuable to investigate necessary conditions to check whether our bounds are tight. Furthermore, extending our analysis to lower-dimensional reward classes could reduce the complexity of learning transferable rewards. Although our paper focuses on discrete state and action spaces, an exciting avenue for future research would be to extend our results to continuous state and action spaces, which are more commonly encountered in practice. Finally, as our work is mainly theoretical, experimental validation on real-world applications could provide valuable insight into the practical aspects and challenges of transferability.
@@ Appendix A Notations
$\ell(r,\mu^{\prime})$ :$\vcentcolon=\quad\max_{\mu\in\mathcal{M}}J(r,\mu)-J(r,\mu^{\prime})$
@@ Appendix B Conjugate duality in regularized IRL
For any proper convex function $f^{*}:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ it holds
@@ Appendix C Regularizers
In the limit $\alpha\to 1$, the Tsallis entropy equals the Shannon entropy defined above. However, in this paper, we use Tsallis entropy to refer to the choice $\alpha=1/2$, which is often adopted as regularization in multi-armed bandit and, more recently, policy optimization algorithms [Zimmert and Seldin, 2021, Dann et al., 2023]. In particular, we consider $h(\pi_{s})=-\tau\mathcal{H}_{1/2}(\pi_{s})=-2\tau\left({\sum_{a}\sqrt{\pi_{s}(a)}-1}\right)$ for some $\tau>0$. We have $0\leq-h\leq 2\tau\left({\sqrt{|\mathcal{A}|}-1}\right)$ and it can be shown that the optimal policy satisfies
@@ D.1 Lipschitz continuity from policies to occupancy measures
For any $\mu_{1},\mu_{2}\in\mathcal{M}$, we have
@@
where in the last step we used
$\left\lVert\nu_{1}-\nu_{2}\right\rVert_{1}$ $=\gamma\sum_{s}\left|{\sum_{s^{\prime},a^{\prime}}P(s|s^{\prime},a^{\prime})\left({\mu_{1}(s^{\prime},a^{\prime})-\mu_{2}(s^{\prime},a^{\prime})}\right)}\right|$ (27)
$\leq\gamma\sum_{s^{\prime},a^{\prime}}\underbrace{\sum_{s}P(s|s^{\prime},a^{\prime})}_{=1}\left|{\left({\mu_{1}(s^{\prime},a^{\prime})-\mu_{2}(s^{\prime},a^{\prime})}\right)}\right|$ (28)
$=\gamma\left\lVert\mu_{1}-\mu_{2}\right\rVert_{1}.$ (29)
By rearranging terms we arrive at the desired inequality
@@ D.2 Strong convexity
The strong convexity part of Proposition D.5 follows from the following more general result.
@@
Let Assumption 2 hold. Suppose that $h$ is $\eta_{h}$-strongly convex with respect to the $\left\lVert\cdot\right\rVert_{1}$ norm. Then, $\bar{h}$ is $\eta_{h}\nu_{\text{min}}/H_{\gamma}^{2}$-strongly convex with respect to $\left\lVert\cdot\right\rVert_{1}$.
@@
We need to show that for $\alpha\in(0,1)$ and $\bar{\alpha}=1-\alpha$ and any two $\mu_{1},\mu_{2}\in\mathcal{M}$ with $\mu_{1}\neq\mu_{2}$ it holds that
@@ D.3 Lipschitz gradients
In this section, we show how we can get bounds on the Lipschitz constant $L_{\mathcal{K}}$ from Assumption 3.3. To this end, we first need to lower bound the optimal policies.
@@
In the following, we establish upper bounds for the Hessians of the occupancy measure regularizations, $\bar{h}$, resulting from Shannon and Tsallis entropy regularization of the policy. In particular, we aim to upper bound the maximum norm of the Hessian in terms of the smallest entry of the policy.
@@
Consider $\mu\in\mathcal{M}$ and let $\pi_{\text{min}}=\min_{s,a}\pi^{\mu}(a|s)$. Then, the Hessian of $\bar{h}$ is upper bounded as follows:
@@ D.4 Dual smoothness and strong convexity
    Let $H_{\gamma}\vcentcolon=1/(1-\gamma)$, $R\vcentcolon=\max_{r\in\mathcal{R}}\left\lVert r\right\rVert_{\infty}$, $D=\max_{r,r^{\prime}\in\mathcal{R}}\left\lVert r-r^{\prime}\right\rVert_{2}$, and suppose that $\tau<2D$, then for the Shannon entropy the inequality (50) holds with
    $\sigma_{\mathcal{R}}=\dfrac{\nu_{\text{min}}\left({1-\frac{\tau}{2D}}\right)\exp\left({\frac{-2RH_{\gamma}}{\tau}}\right)}{D|\mathcal{S}||\mathcal{A}|^{2+H_{\gamma}}},$ (51)
@@
    $\sigma_{\mathcal{R}}=\dfrac{\nu_{\text{min}}\left({1-\frac{\tau}{2D}}\right)}{\sqrt{2}\left({\left({2R/\tau+3\sqrt{|\mathcal{A}|}}\right)H_{\gamma}}\right)^{3}D|\mathcal{S}||\mathcal{A}|}.$ (52)
@@
=$\stackrel{{\scriptstyle(ii)}}{{=}}\sup_{\bar{\mu}\in\mathcal{K}_{\epsilon}}\left[{\langle r^{\prime},\bar{\mu}\rangle-f(\mu)-\langle r,\bar{\mu}-\mu\rangle-\dfrac{L}{2}\left\lVert\bar{\mu}-\mu\right\rVert_{2}^{2}}\right]$
@@
Here, $(i)$ follows from $\mathcal{K}_{\epsilon}\subset\mathcal{M}$, $(ii)$ from the fact that Lipschitz gradients imply that
$f(\bar{\mu})\geq f(\mu)+\langle g,\bar{\mu}-\mu\rangle+\dfrac{L}{2}\left\lVert\bar{\mu}-\mu\right\rVert_{2}^{2},$ (54)
@@
respectively. Furthermore, for any $\mu\in\mathcal{K}_{\epsilon}$, we have by Proposition D.1 the lower bound
@@
which is larger than zero as long as $\tau<2D$. Plugging in the Lipschitz constants for the two regularizations yields
$\sigma_{\mathcal{R}}=\dfrac{2\pi_{\text{min, Sh}}^{\prime}\nu_{\text{min}}}{D|\mathcal{S}||\mathcal{A}|}\left({1-\dfrac{\tau}{2D}}\right)=\dfrac{\pi_{\text{min, Sh}}\nu_{\text{min}}}{D|\mathcal{S}||\mathcal{A}|}\left({1-\dfrac{\tau}{2D}}\right)=\dfrac{\exp\left({-2RH_{\gamma}/\tau}\right)\nu_{\text{min}}}{D|\mathcal{S}||\mathcal{A}|^{2+H_{\gamma}}}\left({1-\dfrac{\tau}{2D}}\right),$ (62)
@@
$\sigma_{\mathcal{R}}$ $=\dfrac{2{\pi_{\text{min, Ts}}^{\prime}}^{3/2}\nu_{\text{min}}}{D|\mathcal{S}||\mathcal{A}|}\left({1-\dfrac{\tau}{2D}}\right)=\dfrac{\pi_{\text{min, Ts}}^{3/2}\nu_{\text{min}}}{\sqrt{2}D|\mathcal{S}||\mathcal{A}|}\left({1-\dfrac{\tau}{2D}}\right)$ (63)
$=\dfrac{\nu_{\text{min}}}{\sqrt{2}\left({\left({2R/\tau+3\sqrt{|\mathcal{A}|}}\right)H_{\gamma}}\right)^{3}D|\mathcal{S}||\mathcal{A}|}\left({1-\dfrac{\tau}{2D}}\right),$ (64)
@@ D.5 Regularity constants
In the following Proposition, we summarize the regularity constants for Shannon and Tsallis entropy regularization. We highlight that these constants are lower bounds for $\eta$ and $\sigma_{\mathcal{R}}$. {restatable}propositionstrongconvexitylipschitzgradients Let $H_{\gamma}\vcentcolon=1/(1-\gamma)$, $R\vcentcolon=\max_{r\in\mathcal{R}}\left\lVert r\right\rVert_{\infty}$, and $D=\max_{r,r^{\prime}\in\mathcal{R}}\left\lVert r-r^{\prime}\right\rVert_{2}$. Suppose that $\tau<2D$, then for the Shannon entropy, Assumption 3.3 holds with
$\eta=\tau\nu_{\text{min}}/H_{\gamma}^{2}\quad\text{and}\quad\sigma_{\mathcal{R}}=\dfrac{\nu_{\text{min}}\left({1-\frac{\tau}{2D}}\right)\exp\left({\frac{-2RH_{\gamma}}{\tau}}\right)}{D|\mathcal{S}||\mathcal{A}|^{2+H_{\gamma}}},$ (65)
@@
$\eta=\tau\nu_{\text{min}}/(2H_{\gamma}^{2}|\mathcal{A}|)\quad\text{and}\quad\sigma_{\mathcal{R}}=\dfrac{\nu_{\text{min}}\left({1-\frac{\tau}{2D}}\right)}{\sqrt{2}\left({\left({2R/\tau+3\sqrt{|\mathcal{A}|}}\right)H_{\gamma}}\right)^{3}D|\mathcal{S}||\mathcal{A}|}.$ (66)
@@ D.6 Suboptimality bounds
$\ell(r,\mu^{\prime})$ $=\langle r,\mu-\mu^{\prime}\rangle-\bar{h}(\mu)+\bar{h}(\mu^{\prime})$ (67)
@@
$\ell(r,\mu^{\prime})\leq 2\left\lVert[r]_{\mathcal{U}}-[r^{\prime}]_{\mathcal{U}}\right\rVert_{\infty}\leq 2\left\lVert[r]_{\mathcal{U}}-[r^{\prime}]_{\mathcal{U}}\right\rVert_{2}.$ (70)
@@
$\ell(r,\mu^{\prime})$ $=\max_{\mu\in\mathcal{M}}J(r,\mu)-J(r,\mu^{\prime})$
@@ D.7 Perturbation bounds
Example E.1.
We consider a two-state, two-action MDP with $\mathcal{S}=\mathcal{A}=\left\{{0,1}\right\}$, uniform initial state distribution, discount rate $\gamma=0.9$, and Shannon entropy regularization $h=-\mathcal{H}$ (see Appendix C). Suppose the expert reward is $r^{\textsf{E}}(s,a)=\mathbbm{1}\{s=1\}$ and consider the transition laws, $P^{0}$ and $P^{1}$, defined by $P^{0}(0|s,a)=1$ and $P^{1}(0|s,a)=\beta\cdot\mathbbm{1}\left\{{s=0,a=0}\right\}$ for some $\beta\in(0,1)$. Also, consider the two experts $\mu^{\textsf{E}}_{P^{0}}=\mathsf{RL}_{P^{0}}(r^{\textsf{E}})$ and $\mu^{\textsf{E}}_{P^{1}}=\mathsf{RL}_{P^{1}}(r^{\textsf{E}})$, and suppose we recovered the reward $\hat{r}(s,a)=-r^{\textsf{E}}$. Then, the following holds: 1) We have $\ell_{P^{0}}(\hat{r},\mu^{\textsf{E}}_{P^{0}})=0$ and $\ell_{P^{1}}(\hat{r},\mu^{\textsf{E}}_{P^{1}})=\mathcal{O}(\beta)$. That is, for small $\beta$, the reward $\hat{r}$ is a good solution to the IRL problem, as both experts are approximately optimal under $\hat{r}$. 2) The rank condition (9) between $P^{0}$ and $P^{1}$ is satisfied for any $\beta\in(0,1)$. 3) For a new transition law $P$ defined by $P(0|s,a)=\mathbbm{1}\left\{{s=1,a=0}\right\}$, we have $\ell_{P}(r^{\textsf{E}},\mathsf{RL}_{P}(\hat{r}))\approx 4.81$, i.e. $\mathsf{RL}_{P}(\hat{r})$ performs poorly under the experts’ reward.
In the following we prove the claims 1. and 2., while 3. is computed via regularized dynamic programming [Geist et al., 2019]. [^2] The code is provided as a .zip file in the supplementary material.
@@
    Consider the transition law $P^{\prime}$ defined by $P^{\prime}(0|s,a)=0$. Observe that while $\left\lVert P^{0}-P^{\prime}\right\rVert$ is large, the potential shaping spaces $\mathcal{U}_{P^{0}}$ and $\mathcal{U}_{P^{\prime}}$ coincide. Moreover, we have
@@
    In light of Propositions D.6, D.7, and 3.4, we have
    $\ell_{P^{1}}(\hat{r},\mathsf{RL}_{P^{1}}(r^{\textsf{E}}))$ $\leq 2\left\lVert[\hat{r}]_{\mathcal{U}_{P^{1}}}-[r^{\textsf{E}}]_{\mathcal{U}_{P^{1}}}\right\rVert_{2}$ (83)
@@
    $\leq 2\sqrt{2}\gamma\cdot H_{\gamma}\cdot\left\lVert P-P^{\prime}\right\rVert\leq 4\gamma\cdot H_{\gamma}\cdot\beta.$ (85)
@@
    To this end, we choose the matrix representation
    $E=\begin{bmatrix}I\\ I\end{bmatrix}\quad\text{and}\quad P=\begin{bmatrix}P^{0}\\ P^{1}\end{bmatrix},$ (87)
    where $I\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$ is the identity matrix and $P^{0},P^{1}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$ are the state transition matrices corresponding to the actions $0,1$, respectively. Let $C=\begin{bmatrix}E-\gamma P^{0},&&E-\gamma P^{1}\end{bmatrix}$. We have
    $P^{0}=\begin{bmatrix}1&0\\ 1&0\\ 1&0\\ 1&0\end{bmatrix}\quad\text{and}\quad P^{1}=\begin{bmatrix}\beta&1-\beta\\ 0&1\\ 0&1\\ 0&1\end{bmatrix},$ (88)

### v2

@@ 5 Experiments
To validate our results experimentally, we adopt a stochastic variant of the WindyGridworld environment (Sutton and Barto, 2018). In this environment, the agent moves to the intended grid cell with a probability of $(1-\beta)$ and is pushed one step further in the direction of the wind with a probability of $\beta$. Using Algorithm 1, we recover a reward $\hat{r}$ from demonstrations of two experts, both exposed to the same wind strength $\beta$ but different wind directions – North and East. The experiments are repeated for a varying number of expert demonstrations $N^{\textsf{E}}\in\{10^{3},10^{4},10^{5},10^{6}\}$ and wind strengths $\beta\in\{0.01,0.1,0.5,1.0\}$. We then test the transferability to two different environments: one with South wind, $P^{\text{South}}$, and a zero-wind environment with cyclically shifted actions, $P^{\text{Shifted}}$. Figure 2(a) shows that the second principal angle between the two experts’ transition laws $P_{0}$ and $P_{1}$ increases with increasing wind strength. Moreover, Figure 2(b)-(d) show that both the closeness between $\hat{r}$ and $r^{\textsf{E}}$ in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$ and the transferability to $P^{\text{South}}$ and $P^{\text{Shifted}}$ improve with a larger second principal angle, as expected from Theorem 3.4. For a more detailed discussion of the experiments we refer to Appendix K.
@@ 6 Conclusion
Our results provide only sufficient conditions for transferability. It would be valuable to investigate necessary conditions to check whether our bounds are tight. Furthermore, extending our analysis to lower-dimensional reward classes could reduce the complexity of learning transferable rewards. Although our paper focuses on discrete state and action spaces, an exciting avenue for future research would be to extend our results to continuous state and action spaces, which are more commonly encountered in practice. We expect that our proof methods can be generalized to this setting, but the analysis will be more intricate due to the infinite-dimensional reward and occupancy measure spaces. Finally, as our work is mainly theoretical, experimental validation on real-world applications could provide valuable insight into the practical aspects and challenges of transferability.
@@ Appendix A Notations
$\mathsf{SubOpt}(r,\mu^{\prime})$ :$\vcentcolon=\quad\max_{\mu\in\mathcal{M}}J(r,\mu)-J(r,\mu^{\prime})$
@@ Appendix B Conjugate duality in regularized IRL
For any proper convex function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ it holds
@@ Appendix C Regularizers
In the limit $\alpha\to 1$, the Tsallis entropy equals the Shannon entropy defined above. However, in this paper, we use Tsallis entropy to refer to the choice $\alpha=1/2$, which is often adopted as regularization in multi-armed bandit and, more recently, policy optimization algorithms [Zimmert and Seldin, 2021, Dann et al., 2023]. That is, we consider $h(\pi_{s})=-\tau\mathcal{H}_{1/2}(\pi_{s})=-2\tau\left({\sum_{a}\sqrt{\pi_{s}(a)}-1}\right)$ for some $\tau>0$. We have $0\leq-h\leq 2\tau\left({\sqrt{|\mathcal{A}|}-1}\right)$ and it can be shown that the optimal policy satisfies
@@ D.1 Lipschitz continuity from policies to occupancy measures
Let Assumption 2 hold. For any $\mu_{1},\mu_{2}\in\mathcal{M}$, we have
@@
where in the last step we used again that $\left\lVert\nu_{1}-\nu_{2}\right\rVert_{1}\leq\gamma\left\lVert\mu_{1}-\mu_{2}\right\rVert_{1}$. By rearranging terms we arrive at the desired inequality
@@ D.2 Strong convexity
Next, we show that strong convexity of the policy regularizer translates into strong convexity in the occupancy measure.
@@
Let Assumption 2 hold and suppose that $h$ is $\eta_{h}$-strongly convex with respect to the $\left\lVert\cdot\right\rVert_{1}$ norm. Then, $\bar{h}$ is $\eta_{h}\nu_{\text{min}}/H_{\gamma}^{2}$-strongly convex with respect to $\left\lVert\cdot\right\rVert_{1}$.
@@
We need to show that for $\alpha\in(0,1),\bar{\alpha}=1-\alpha$ and any two $\mu_{1},\mu_{2}\in\mathcal{M}$ it holds that
@@ D.3 Lipschitz gradients
In this section, we show how we can get bounds on the Lipschitz constant $L_{\mathcal{K}}$. To this end, we first need to lower bound the optimal policies.
@@
In the following, we establish upper bounds for the Hessians of the occupancy measure regularizations, $\bar{h}$, resulting from Shannon and Tsallis entropy regularization of the policy. In particular, we aim to upper-bound the maximum norm of the Hessian in terms of the smallest entry of the policy.
@@
Let Assumption 2 hold. Consider $\mu\in\mathcal{M}$ and let $\pi_{\text{min}}=\min_{s,a}\pi^{\mu}(a|s)>0$. Then, the Hessian of $\bar{h}$ is upper bounded as follows:
@@ D.4 Dual smoothness and strong convexity
    Let $H_{\gamma}\vcentcolon=1/(1-\gamma)$, $R\vcentcolon=\max_{r\in\mathcal{R}}\left\lVert r\right\rVert_{\infty}$, $D=\max_{r,r^{\prime}\in\mathcal{R}}\left\lVert r-r^{\prime}\right\rVert_{2}$, and suppose that $\tau<D$, then for the Shannon entropy the inequality (47) holds with
    $\sigma_{\mathcal{R}}=\dfrac{\nu_{\text{min}}\exp\left({\frac{-2RH_{\gamma}}{\tau}}\right)}{2D|\mathcal{S}||\mathcal{A}|^{2+H_{\gamma}}},$ (48)
@@
    $\sigma_{\mathcal{R}}=\dfrac{\nu_{\text{min}}}{2\sqrt{2}D|\mathcal{S}||\mathcal{A}|\left({\left({2R/\tau+3\sqrt{|\mathcal{A}|}}\right)H_{\gamma}}\right)^{3}}.$ (49)
@@
$\stackrel{{\scriptstyle(ii)}}{{\geq}}\sup_{\bar{\mu}\in\mathcal{K}_{\epsilon}}\left[{\langle r^{\prime},\bar{\mu}\rangle-f(\mu)-\langle r,\bar{\mu}-\mu\rangle-\dfrac{L}{2}\left\lVert\bar{\mu}-\mu\right\rVert_{2}^{2}}\right]$
@@
Here, $(i)$ follows from $\mathcal{K}_{\epsilon}\subset\mathbb{R}^{\mathcal{S}\times\mathcal{A}}$, $(ii)$ from the fact that Lipschitz gradients imply that
$f(\bar{\mu})\leq f(\mu)+\langle g,\bar{\mu}-\mu\rangle+\dfrac{L}{2}\left\lVert\bar{\mu}-\mu\right\rVert_{2}^{2},$ (51)
@@
respectively. Furthermore, for any $\mu\in\mathcal{K}_{\epsilon}$, we have by Proposition D.1 and equivalence of norms
@@
which is larger than $\tau/LD$ for $\tau<D$. We can therefore choose $\sigma_{\mathcal{R}}=\tau/LD$ as a dual strong convexity constant. Plugging in the Lipschitz constants for the two regularizations yields
$\sigma_{\mathcal{R},\text{Sh}}=\dfrac{\pi_{\text{min, Sh}}^{\prime}\nu_{\text{min}}}{D|\mathcal{S}||\mathcal{A}|}=\dfrac{\pi_{\text{min, Sh}}\nu_{\text{min}}}{2D|\mathcal{S}||\mathcal{A}|}=\dfrac{\exp\left({-2RH_{\gamma}/\tau}\right)\nu_{\text{min}}}{2D|\mathcal{S}||\mathcal{A}|^{2+H_{\gamma}}},$ (59)
@@
$\sigma_{\mathcal{R},\text{Ts}}$ $=\dfrac{{\pi_{\text{min, Ts}}^{\prime}}^{3/2}\nu_{\text{min}}}{D|\mathcal{S}||\mathcal{A}|}=\dfrac{\pi_{\text{min, Ts}}^{3/2}\nu_{\text{min}}}{2\sqrt{2}D|\mathcal{S}||\mathcal{A}|}=\dfrac{\nu_{\text{min}}}{2\sqrt{2}D|\mathcal{S}||\mathcal{A}|\left({\left({2R/\tau+3\sqrt{|\mathcal{A}|}}\right)H_{\gamma}}\right)^{3}},$ (60)
@@
Remark D.6 (Large $\tau$ regime).
Note that if $\tau\geq 2D/\sqrt{|\mathcal{S}||\mathcal{A}|}$, we can set $\sigma_{\mathcal{R}}=\alpha=1/L$, while still satisfying the condition (57). This leads to the strong convexity constants
$\sigma_{\mathcal{R},\text{Sh}}=\dfrac{\exp\left({-2RH_{\gamma}/\tau}\right)\nu_{\text{min}}}{2\tau|\mathcal{S}||\mathcal{A}|^{2+H_{\gamma}}},\;\sigma_{\mathcal{R},\text{Ts}}=\dfrac{\nu_{\text{min}}}{2\sqrt{2}\tau|\mathcal{S}||\mathcal{A}|\left({\left({2R/\tau+3\sqrt{|\mathcal{A}|}}\right)H_{\gamma}}\right)^{3}}.$
@@ D.5 Regularity constants
In the following Proposition, we summarize the regularity constants for Shannon and Tsallis entropy regularization. We highlight that these constants are lower bounds for $\eta$ and $\sigma_{\mathcal{R}}$. {restatable}propositionstrongconvexitylipschitzgradients Let $H_{\gamma}\vcentcolon=1/(1-\gamma)$, $R\vcentcolon=\max_{r\in\mathcal{R}}\left\lVert r\right\rVert_{\infty}$, and $D=\max_{r,r^{\prime}\in\mathcal{R}}\left\lVert r-r^{\prime}\right\rVert_{2}$. Suppose that $\tau<D$, then for the Shannon entropy, Assumption 3.3 holds with
$\eta=\tau\nu_{\text{min}}/H_{\gamma}^{2}\quad\text{and}\quad\sigma_{\mathcal{R}}=\dfrac{\exp\left({-2RH_{\gamma}/\tau}\right)\nu_{\text{min}}}{2D|\mathcal{S}||\mathcal{A}|^{2+H_{\gamma}}},$ (61)
@@
$\eta=\tau\nu_{\text{min}}/(2H_{\gamma}^{2}|\mathcal{A}|)\quad\text{and}\quad\sigma_{\mathcal{R}}=\dfrac{\nu_{\text{min}}}{2\sqrt{2}D|\mathcal{S}||\mathcal{A}|\left({\left({2R/\tau+3\sqrt{|\mathcal{A}|}}\right)H_{\gamma}}\right)^{3}}.$ (62)
@@ D.6 Suboptimality bounds
$\mathsf{SubOpt}(r,\mu^{\prime})$ $=\langle r,\mu-\mu^{\prime}\rangle-\bar{h}(\mu)+\bar{h}(\mu^{\prime})$ (63)
@@
$\mathsf{SubOpt}(r,\mu^{\prime})\leq 2\left\lVert[r]_{\mathcal{U}}-[r^{\prime}]_{\mathcal{U}}\right\rVert_{\infty}\leq 2\left\lVert[r]_{\mathcal{U}}-[r^{\prime}]_{\mathcal{U}}\right\rVert_{2}.$ (66)
@@
$\mathsf{SubOpt}(r,\mu^{\prime})$ $=\max_{\mu\in\mathcal{M}}J(r,\mu)-J(r,\mu^{\prime})$
@@ D.7 Perturbation bounds
Example 3.2. We consider a two-state, two-action MDP with $\mathcal{S}=\mathcal{A}=\left\{{0,1}\right\}$, uniform initial state distribution, discount rate $\gamma=0.9$, and Shannon entropy regularization $h=-\mathcal{H}$ (see Appendix C). Suppose the expert reward is $r^{\textsf{E}}(s,a)=\mathbbm{1}\{s=1\}$ and consider the transition laws, $P^{0}$ and $P^{1}$, defined by $P^{0}(0|s,a)=0.75$ and $P^{1}(0|s,a)=0.25+\beta\cdot\mathbbm{1}\left\{{s=0,a=0}\right\}$ for some $\beta\in[0,0.75]$. Also, consider the two experts $\mu^{\textsf{E}}_{P^{0}}=\mathsf{RL}_{P^{0}}(r^{\textsf{E}})$ and $\mu^{\textsf{E}}_{P^{1}}=\mathsf{RL}_{P^{1}}(r^{\textsf{E}})$, and suppose we recovered the reward $\hat{r}(s,a)=-r^{\textsf{E}}$. Then, the following holds: 1) We have $\mathsf{SubOpt}_{P^{0}}(\hat{r},\mu^{\textsf{E}}_{P^{0}})=0$ and $\mathsf{SubOpt}_{P^{1}}(\hat{r},\mu^{\textsf{E}}_{P^{1}})=\mathcal{O}(\beta)$. That is, for small $\beta$, the reward $\hat{r}$ is a good solution to the IRL problem, as both experts are approximately optimal under $\hat{r}$. 2) The rank condition (9) between $P^{0}$ and $P^{1}$ is satisfied for any $\beta>0$. 3) For a new transition law $P$ defined by $P(0|s,a)=\mathbbm{1}\left\{{s=1,a=0}\right\}$, we have $\mathsf{SubOpt}_{P}(r^{\textsf{E}},\mathsf{RL}_{P}(\hat{r}))\approx 4.81$, i.e. $\mathsf{RL}_{P}(\hat{r})$ performs poorly under the experts’ reward.
In the following we prove the claims 1. and 2., while 3. is computed via regularized dynamic programming [Geist et al., 2019]. [^3] The code is openly accessible at https://github.com/andrschl/irl_transferability. Furthermore, we illustrate the occupancy measure spaces corresponding to $P^{0}$ and $P^{1}$ for different $\beta$ in Figure 3.
Figure 3: The set of occupancy measures $\mathcal{M}_{P^{0}}$ and $\mathcal{M}_{P^{1}}$ are illustrated in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}\cong\bm{1}^{\perp}$. For a two-state-two-action MDP, the set of occupancy measures is given by the intersection of a two-dimensional affine subspace (a plane in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$) with the probability simplex in $\mathbb{R}^{4}$ (a tetrahedron in $\mathbb{R}^{\mathcal{S}\times\mathcal{A}}/\bm{1}$). We see that for a small $\beta$, the sets $\mathcal{M}_{P^{0}}$ and $\mathcal{M}_{P^{1}}$ are approximately parallel. That is, the angle between their normal vectors, which span the potential shaping spaces $\mathcal{U}_{P^{0}}$ and $\mathcal{U}_{P^{1}}$, is small. In contrast, for a large $\beta$ the orientation of $\mathcal{M}_{P^{0}}$ and $\mathcal{M}_{P^{1}}$ is very different, resulting in a large angle between the corresponding normal vectors.
@@
    Consider the transition law $P^{\prime}$ defined by $P^{\prime}(0|s,a)=0.25$. First, we observe that while $\left\lVert P^{0}-P^{\prime}\right\rVert$ is large, the potential shaping spaces $\mathcal{U}_{P^{0}}$ and $\mathcal{U}_{P^{\prime}}$ coincide. To see this note that we have $P^{\prime}(\cdot|s,a)=P^{0}(\cdot|s,a)+\Delta$, where $\Delta=[-0.5,0.5]^{\top}$. Hence, we have for any $x\in\mathbb{R}^{2}$ that
    $(E-\gamma P^{\prime})x=(E-\gamma P^{0})x-\gamma\langle\Delta,x\rangle\bm{1}_{4}=(E-\gamma P^{0})\left({x-\frac{\gamma\langle\Delta,x\rangle}{1-\gamma}\bm{1}_{2}}\right),$
    where $\bm{1}_{n}$ denotes the all-one vector in $\mathbb{R}^{n}$. Therefore, $\operatorname{\operatorname{span}}(E-\gamma P^{\prime})=\operatorname{\operatorname{span}}(E-\gamma P^{0})$. Moreover, we have
@@
    In light of Propositions D.7, D.8, and 3.4, this implies that
    $\mathsf{SubOpt}_{P^{1}}(\hat{r},\mathsf{RL}_{P^{1}}(r^{\textsf{E}}))$ $\leq 2\left\lVert[\hat{r}]_{\mathcal{U}_{P^{1}}}-[r^{\textsf{E}}]_{\mathcal{U}_{P^{1}}}\right\rVert_{2}$ (79)
@@
    $\leq 2\gamma\cdot H_{\gamma}\cdot\left\lVert P-P^{\prime}\right\rVert\leq 2\sqrt{2}\gamma\cdot H_{\gamma}\cdot\beta.$ (81)
@@
    By the same reasoning as above, we can equivalently show the rank condition for the transition laws ${P^{0}}^{\prime},{P^{1}}^{\prime}$ defined by ${P^{0}}^{\prime}(0|s,a)=1$ and ${P^{1}}^{\prime}(0|s,a)=\beta\cdot\mathbbm{1}\left\{{s=0,a=0}\right\}$. To this end, we choose the matrix representation
    $E=\begin{bmatrix}I\\ I\end{bmatrix}\quad\text{and}\quad P=\begin{bmatrix}P_{a_{0}}\\ P_{a_{1}}\end{bmatrix},$ (83)
    where $I\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$ is the identity matrix and $P_{a_{0}},P_{a_{1}}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}|}$ are the state transition matrices corresponding to the actions $0,1$, respectively. Let $C=\begin{bmatrix}E-\gamma{P^{0}}^{\prime},&&E-\gamma{P^{1}}^{\prime}\end{bmatrix}$. We have
    ${P^{0}}^{\prime}=\begin{bmatrix}1&0\\ 1&0\\ 1&0\\ 1&0\end{bmatrix}\quad\text{and}\quad{P^{1}}^{\prime}=\begin{bmatrix}\beta&1-\beta\\ 0&1\\ 0&1\\ 0&1\end{bmatrix},$ (84)